OpenStudy (anonymous):

Place the following in increasing order: 3/5 , 70%, 0.8, 11/20 , 65%, 0.09.

OpenStudy (anonymous):

WILL MEDAL

OpenStudy (anonymous):

oh ok so which one is the smallest

OpenStudy (anonymous):

hello? if you need more help you can just ask :]

OpenStudy (lichking):

should i post answer or help you to solve?

OpenStudy (anonymous):

we shouldn't just give the answer tho :(

OpenStudy (lichking):

ok not here but i can pm the answer...there are no rules against that! XD

OpenStudy (anonymous):

lol kk

OpenStudy (anonymous):

ok forget it im just gonna explain the answer so that i give the answer but also help her

OpenStudy (anonymous):

so first find the smallest which is the decimal [O.8 or O.O9 i will let you figure that out] then you find the other smallest... then you will get the answer

OpenStudy (anonymous):

i dont understand can you tell me the answer @LichKing

OpenStudy (sloppycanada):

Alrighty then! Let's start by making our life super easy and making them all the same format. Personally, I like decimals, but your choice.

OpenStudy (sloppycanada):

3/5 = 60% = .60 70% = 7/10 = .70 .8 = .80 11/20 = .55 65% = .65 .09 = .09 Now out of all of these, since they're the same format, which is smallest?

OpenStudy (anonymous):

the smallest is .o9 ?

OpenStudy (anonymous):

@sloppycanada

OpenStudy (sloppycanada):

Yup, next?

OpenStudy (anonymous):

wow you must have explained it better then me :D good job @sloppycanada